physics

A 75-kg swimmer dives horizontally off a 500-kg raft. If the diver's speed immediately after leaving
the raft is 4 m/s, what is the corresponding raft speed?
Answer #1
75 * v + 500 * 4 = 0
v = 26 m /s
answered by: cayla
Answer #2
Conservation of momentum

75*4 = 500*x

x = .6
answered by: slim stud
Answer #3
75 * 4+ 500 * v = 0
v = 0.6 m /s
answered by: dfsfd